OpenStudy (anonymous):

4-c/2c-8 Please show all work, and state any excuded values

Parth (parthkohli):

\[4 - c = -(c - 4)\]Right?

OpenStudy (anonymous):

idk..

OpenStudy (anonymous):

is it like this \[\Large \frac{ 4-c }{ 2c }-8\] ??

OpenStudy (anonymous):

No like this @kausarsalley \[\frac{ 4-c }{ 2c-8 }\]

OpenStudy (jhannybean):

factor out a 2 at the bottom...so you'd have (4- c)/2(c - 4)

OpenStudy (anonymous):

so like \[\frac{ 4c }{ 2(c-4) }\]

OpenStudy (jhannybean):

now you have to find the zeroes?

OpenStudy (anonymous):

Hmm?

OpenStudy (jhannybean):

no it would look like \[\large \frac{4-c}{2(c-4)}\]

OpenStudy (anonymous):

ooh oops i forgot the minus

OpenStudy (anonymous):

ok now what?

OpenStudy (jhannybean):

for excluded values, equal the numerator and denominator to 0 and solve for c.

OpenStudy (anonymous):

Ok factoring, right?

Parth (parthkohli):

\[\dfrac{4 - c}{2(c - 4)}\]can be simplified

OpenStudy (anonymous):

Oh yea...distribute

OpenStudy (anonymous):

so \[\frac{ 4-c }{ 2c-8 }\]

OpenStudy (anonymous):

?

Parth (parthkohli):

Uh no.\[\dfrac{4 - c}{2(c -4)} = \dfrac{-(c - 4)}{2(c - 4)}\]

OpenStudy (jhannybean):

well the negative......

OpenStudy (anonymous):

Ok..why don't you distribute..?

Parth (parthkohli):

Why would we distribute? That just gets us back to the original question lol

OpenStudy (anonymous):

oh yea

OpenStudy (jhannybean):

Because like he showed above, those can cancel out, leaving you with -1 ontop.

OpenStudy (anonymous):

so what we are doing is un-distributing...kinda?

OpenStudy (jhannybean):

we've already un-distributed haha

OpenStudy (jhannybean):

we're....re-simplifying.

Parth (parthkohli):

We're factorizing.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so after -(c-4)/2(c-4) what would it be?

OpenStudy (jhannybean):

what does (c-4)/(c-4) = ?

OpenStudy (anonymous):

they cancel out

OpenStudy (jhannybean):

so what's left over?

OpenStudy (anonymous):

so you are left with 2 does not equal -1?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

:D was i right?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

\[\large -\frac{\cancel{(c-4)}}{2\cancel{(c-4)}}=-\frac{1}{2}\]

OpenStudy (anonymous):

Ohhh that makes so much sense :D I love you (lol like a sister )

OpenStudy (anonymous):

ok what is theexcuded value

OpenStudy (anonymous):

is it c doesnt equal -1/2? @jhannybean

OpenStudy (anonymous):

\[c \neq-\frac{ 1 }{ }\]

OpenStudy (anonymous):

\[c \neq-\frac{ 1 }{ 2 }\]***

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@jh

OpenStudy (anonymous):

@Jhannybean ***

OpenStudy (jhannybean):

Hm... not too sure.

OpenStudy (anonymous):

actually its \[c \neq4\]